OpenStudy (anonymous):

solve and check 2^x = 4^x + 1

OpenStudy (anonymous):

=fales

OpenStudy (anonymous):

\[ x=-\frac{i \pi }{3 \ln (2)},\quad x=\frac{i \pi }{3 \ln (2)} \]

OpenStudy (anonymous):

The options are -1, -2, 2, or 0.

OpenStudy (lgbasallote):

im guessing that's \( 4^{x + 1}\)

OpenStudy (lgbasallote):

so it would be \(\Large 2^x = 4^{x+1} \rightarrow 2^x = 2^{2(x+1)}\) any ideas now?

OpenStudy (anonymous):

None of -1, -2, 2, or 0. is a solution if your equation is \[ 2^x = 4^x +1 \]

OpenStudy (australopithecus):

oh oh it is 2^(x) = 4^(x+1) xln(2) = (x+1)ln(4) xln(2) = xln(4) + ln(4) xln(2) - xln(4) = ln(4) x(ln(2) - ln(4)) = ln(4) x = ln(4)/(ln(2) - ln(4) x = -2

OpenStudy (australopithecus):

if you need to me to explain my method I can :)

OpenStudy (australopithecus):

@lgbasallote thanks for pointing out the correct form of the question

OpenStudy (lgbasallote):

another method: 2^x = 2^(2x +2) x = 2x + 2 2x - x = -2 x = -2 same thing :D

OpenStudy (anonymous):

Please put parentheses where they belong. There is a big difference between 4^x+1 and 4^(x+1)